Gold forms a substitutional solid solution with silver. Compute the number of gold atoms per cubic centimeter for a silver-gold alloy that contains 48 wt% Au and 52 wt% Ag. The densities of pure gold and silver are 19.32 and 10.49 g/cm3
Answer #1

density of the resulting substance = 0.48*19.32 + 0.52*10.49

=14.73 g/cm^3

now, mass of 1 cm^3 os the alloy = 14.73 grams

so mass of gold in it = 0.48*14.73

=7.0704 grams

molar mass of Au = 197 g/moles

so number of moles = 7.0704/197

so number of atoms = 6.023*10^23*(7.0704/197)

=2.16*10^22 atoms/cm^3